# Diffusion Trajectory-guided Policy for Long-horizon Robot Manipulation

### v1

@@ IV-B Comparisons with State-of-the-Art Methods
Known Scene Results. This experiment is conducted in the D→D setting, utilizing about 5,000 expert demonstrations for training. The training process takes approximately 1.5 days on 8 NVIDIA 24G RTX 3090 GPUs. As shown in Tab. I, DTP significantly outperforms all baseline methods across all metrics in the context of long-horizon tasks. Specifically, DTP increases the success rate for Task 5 from 0.400 to 0.509 and raises the average successful sequence length from 3.30 to 3.55. Notably, compared to GR-1, our baseline model, DTP enhances performance across all metrics, with the average sequence length increasing by 33.9%. These results indicate that DTP demonstrates superior performance in long-horizon tasks, particularly as the task length increases.
Unseen Scene Results. This experiment is conducted in the ABC→D setting, which is particularly challenging: models are trained using data from environments A, B, and C and then tested in environment D, an unseen setting during the training phase. The training process takes approximately 5 days on 8 NVIDIA 24GB RTX 3090 GPUs. We validated DTP’s generalization capability in a new environment. The results are presented in Tab. I. When compared with the baseline GR-1, there is an increase in the average task completion length from 3.06 to 3.43. Additionally, the success rate for completing Task 5 increased to 0.497, the highest recorded value. Notably, even though our method does not use depth modality for training, it outperformed the 3D Diffuser Actor in these tests. This underscores a critical insight: DTP can effectively guide policy learning for long-horizon robot tasks in unseen settings.
Data Efficiency. Robot data is more costly and scarce compared to vision-language data. To evaluate data efficiency, we trained using only 10% of the full dataset in the ABCD→D setting, randomly selecting around 2,000 expert demonstrations from over 20,000 episodes. Training took approximately 1 day on 8 NVIDIA 24GB RTX 3090 GPUs. The results are shown in Tab. I. While performance declines for all methods compared to training on the full dataset, the best baseline method, GR-1, achieves a success rate of 0.778 with an average length of 2.00. DTP shows clear benefits for long-horizon tasks; as task numbers increase, the success rate rises, and the average length reaches 2.55, outperforming other methods. This highlights DTP’s data efficiency. Imitation learning helps the model learn positional preferences, which are essential in long-horizon tasks. When the robot starts from an unseen position, task failures are more likely. However, DTP guides the robot arm with a diffusion trajectory, providing the correct path. Thus, even with fewer demonstrations, DTP quickly acquires the skills.
@@ IV-C Ablation Studies
In this section, we conduct ablation studies to evaluate how the diffusion trajectory improves policy learning in visual robot manipulation tasks. The diffusion trajectory, our key contribution, significantly boosts the efficiency of imitation policy training by fully utilizing available robot data. Furthermore, when integrated with large-scale pretraining baseline methods, our approach serves as a straightforward and effective enhancement to performance. To measure the effectiveness of our method, we contrast it with two fundamental baselines. The first baseline employs the GR-1 framework (Sec. III-B) without video pretraining, while the second utilizes large-scale video pretraining with the Ego4D dataset [30], also based on GR-1 framework. Two baselines are established to verify the efficacy and compatibility of our method with other approaches.
@@
Diffusion Trajectory Policy Scratch. First, we evaluate our method in the ABC→D and 10% ABCD→D settings, as shown in the upper part of Tab. II. The results demonstrate that our diffusion trajectory method significantly enhances performance even without any pretraining. Specifically, our method not only excels in sequentially completed tasks but also shows notable gains in the average task completion length for long-horizon tasks increase of 23.4%. Notably, the success rate for the task 5, which is indicative of the overall long-horizon success, has risen by 56.6%. When compared with the 3D Diffuser Actor, as shown in Tab. I, despite not utilizing depth modality. This highlights our method’s efficiency and capability in handling complex robot manipulation tasks without the need for depth data.
Diffusion Trajectory Policy with Video Pretrain. As illustrated in the bottom part of Tab. II, the variants utilizing our diffusion trajectory effectively serve as a plugin, boosting baseline model performance to state-of-the-art levels. We evaluated our method under both the ABC→D and 10% ABCD→D settings, and the results consistently show improvements over the traditional scratch training method. This clearly indicates that our approach complements and significantly enhances baseline performance across various benchmarks. Additionally, the success rates for each subsequent task show notable increases, with the growth rate rising from 4.2% in the first task to 23.9% in the fifth task. These outcomes further validate that DTP can substantially improve performance in long-horizon manipulation tasks.
Diffusion Trajectory Model Scaling Law. The last row highlights the initial training stage of our Diffusion Trajectory Model. Increasing the training data allows the model to generate more accurate points, enhancing the DTP. The bottom row demonstrates that even with limited demonstration data for imitation learning, scaling up the training for the diffusion trajectory can significantly improve both the success rate and average task completion length. This experimental setup points to a potential direction: although robot demonstration data is costly to obtain, the data for the DTM is relatively easy to annotate. Individuals only need to sketch a coarse trajectory on an RGB image and associate it with relevant language instructions.
@@ IV-D Real Robot Experiment
Experiment Setup. Our real-world robotic system, depicted in Fig. 4, consists of a Franka Emika Panda robot equipped with three Intel RealSense D435i cameras (left, right, and top views) and a Robotiq gripper. First, We collected 1086 demonstrations for all tasks by teleoperation system [31]. Specifically, we collected 290, 258, 100, 184, and 254 demonstrations for five tasks—PickBread, PickStrawberry, OpenTrash, CloseSideDrawer, and OpenSideDrawer respectively, encompassing object transportation and articulated object manipulation, as illustrated in Fig. 5. Training required approximately 1 day on 4 NVIDIA 24GB RTX3090 GPUs with 20 epochs.
@@
Results. The performance of DTP and baseline methods is summarized in Tab. III. Each task was evaluated over 10 trials, with success rates calculated for comparison. Overall, DTP achieved the highest aggregate success rate across tasks. However, in the PickStrawberry task, DTP underperformed compared to ACT. We attribute this to the small size of the target object, as DTP uses an image input resolution of 224x224, while ACT operates at a higher resolution of 480x640, which likely impacts performance. In long-horizon tasks, the robot arm’s initial pose is determined by the completion of the previous task, resulting in random starting configurations. To evaluate DTP’s robustness in such scenarios, we tested it on the OpenSideDrawer task with randomized initial arm poses. DTP achieved a success rate twice as high as the second-best method. Additionally, in the OpenTrash task, which requires precise alignment to a specific area to open the trash bin, DTP demonstrated superior guidance capabilities. While other baseline methods positioned the arm near the target, they often failed to locate the precise opening mechanism, leading to task failure.
@@
Visualization of Diffusion Trajectory Model. As shown in Fig. 6, we present the overall visualization of the diffusion trajectory generation phase, tested in both the Calvin environment and real-world scenarios. The visualizations demonstrate that the trajectories generated by our diffusion trajectory prediction closely match the ground truth. Even when minor deviations occur, the generated trajectories still align with the robotic arm paths dictated by the language instructions.
| PickBread | PickStrawberry | OpenTrash | CloseSideDrawer | OpenSideDrawer* | |
| ACT | 0.7 | 0.9 | 0.3 | 0.3 | 0.4 |
| BAKU[32] | 0.0 | 0.5 | 0.2 | 0.2 | 0.3 |
| GR1 | 0.7 | 0.7 | 0.2 | 0.4 | 0.4 |
| DTP(Ours) | 0.8 | 0.8 | 0.9 | 0.9 | 0.8 |
@@ V CONCLUSION
The limited availability of robot data poses significant challenges in generalizing long-horizon tasks to unseen robot poses and environments. This paper introduces a diffusion trajectory-guided framework that utilizes diffusion trajectories, generated in the RGB domain, to enhance policy learning in long-horizon robot manipulation tasks. This method facilitates the creation of additional training data through data augmentation or manually crafted labels, thereby generating more accuracy diffusion trajectories. Our approach involves two main stages: first, training a diffusion trajectory model to generate task-relevant trajectories; second, using these trajectories to guide the robot’s manipulation policy. We validated our method through extensive experiments on the CALVIN simulation benchmark, where it outperformed state-of-the-art baselines by an average success rate of 25% across various settings. Our results confirm that our method not only substantially improves performance using only robot data but also effectively complements and enhances baseline performance across various settings in the CALVIN benchmarks. Moreover, our method brings about a remarkable enhancement in the performance of real-world robots.
In future work, we plan to extend our method to other state-of-the-art policies, as we believe that incorporating diffusion trajectories will further enhance their effectiveness. Another potential direction is to obtain the diffusion trajectory label using the camera’s intrinsic and extrinsic parameters, which are not fully available from open-source datasets [33]. Recently, Track-Anything [23] demonstrated strong capabilities in tracking arbitrary objects. We could adopt this method to generate diffusion trajectory labels. Furthermore, with similar tracking methods, we can pretrain on large-scale video datasets to train our diffusion trajectory tasks, similar to video prediction tasks.

### v2

@@ IV-B Comparisons with State-of-the-Art Methods
Results in Seen Environments. In the DD setting, using about 5,000 expert demonstrations, training on 8*3090 GPUs took 1.5 days. As shown in Table I, DTP outperforms all baselines in long-horizon tasks, increasing Task 5’s success rate from 0.400 to 0.509 and the average sequence length from 3.30 to 3.55. Compared to GR-1, DTP boosts performance across all metrics, with a 33.9% increase in average sequence length, demonstrating superior performance as task length increases.
Results in Unseen Environments. In the challenging ABCD setting, models are trained on environments A, B, and C, and tested in the unseen environment D. Training took about 5 days on 8*3090 GPUs. As shown in Table I, DTP demonstrates strong generalization, increasing the average task completion length from 3.06 to 3.43 and achieving a Task 5 success rate of 0.497, the highest recorded. Notably, without using depth modality, DTP outperformed the 3D Diffuser Actor, highlighting its effectiveness in guiding policy learning for long-horizon tasks in unseen settings.
Data Efficiency. In the ABCDD setting, we evaluated data efficiency by training with only 10% of the dataset, using around 2,000 expert demonstrations. Training took about 1 day on 8*3090 GPUs. As shown in Table I, while all methods perform worse with less data, the best baseline, GR-1, achieves a success rate of 0.778 with an average length of 2.00. DTP excels in long-horizon tasks, with a success rate increase and an average length of 2.55, outperforming others. These results demonstrate the data efficiency of DTP. By leveraging diffusion-based trajectories, the policy effectively captures positional preferences that are critical for long-horizon tasks. Furthermore, these 2D-particle trajectories provide guidance to the robot arm, enabling it to acquire skills even with a limited number of demonstrations.
@@ IV-C Ablation Studies
In this section, we conduct ablation studies to assess how diffusion trajectories enhance policy learning in visual robot manipulation tasks. This key contribution significantly improves imitation policy training efficiency by fully utilizing robot data. Integrated with large-scale pretraining baselines, it offers a straightforward performance boost. We compare our method against two baselines: one using the GR-1 framework (Sec. III-B) without video pretraining, and another with large-scale video pretraining using the Ego4D [36] dataset, both based on GR-1. These baselines verify our method’s efficacy and compatibility.
@@
Diffusion Trajectory Policy from Scratch. We evaluate our method in the ABCD and 10% ABCDD settings (see Table II). The diffusion trajectory method significantly enhances performance without pretraining, excelling in sequential tasks and increasing the average task completion length by 23.4%. The success rate for Task 5, indicative of long-horizon success, rises by 56.6%.
Diffusion Trajectory Policy with Video Pretrain. As shown in the bottom part of Table II, our diffusion trajectory variants enhance baseline model performance to state-of-the-art levels. Evaluated in ABCD and 10% ABCDD settings, our method consistently outperforms traditional scratch training, significantly boosting baseline performance. Success rates increase notably, from 4.2% in the first task to 23.9% in the fifth, validating DTP’s effectiveness in long-horizon manipulation tasks.
Diffusion Trajectory Model Scaling Law. The last row highlights the initial training stage of our Diffusion Trajectory Model. Increasing training data improves point accuracy, enhancing DTP. Even with limited demonstration data, scaling diffusion trajectory training boosts success rates and task completion length. This suggests a potential direction: while robot demonstration data is costly, DTM data is easier to annotate, requiring only a coarse trajectory sketch on an RGB image with language instructions.
@@ IV-D Real Robot Experiment
Experiment Setup. Our robotic system features a Franka Emika Panda robot with three Intel RealSense D435i cameras and a Robotiq gripper. We collected 1,512 demonstrations using a teleoperation system [37], with 290, 258, 100, 184, and 254 demonstrations for the tasks PickBread, PickStrawberry, OpenTrash, CloseSideDrawer, and OpenSideDrawer, respectively, covering object transportation and manipulation (see Fig. 4a). Additionally, we collected 426 demonstrations for a long-horizon task comprising three subtasks(A-B-C) (see Fig. 4b). In real robot setup, the number of generated actions is set to 25. This adjustment addresses sim-to-real discrepancies and latency, as setting both values to 1 causes the robot to remain almost completely still, with only minor vibrations. The number of executed actions remains at 1. We also incorporate an action ensemble mechanism, similar to ALOHA [38]. The training process spanned 20 epochs and took approximately one day on four RTX 3090 GPUs.
@@
Results. The performance of DTP and baseline methods in five individual tasks is summarized in Table III upper part. Each task was evaluated over 10 trials, with success rates calculated for comparison. Overall, DTP achieved the highest aggregate success rate across tasks. However, in the PickStrawberry task, DTP underperformed compared to ACT. We attribute this to the small size of the target object, as DTP uses an image input resolution of 224x224, while ACT operates at a higher resolution of 480x640, which likely impacts performance. In long-horizon tasks, the robot arm’s initial pose is determined by the completion of the previous task, resulting in random starting configurations. To evaluate DTP’s robustness in such scenarios, we tested it on the OpenSideDrawer task with randomized initial arm poses. DTP achieved a success rate twice as high as the second-best method. Additionally, in the OpenTrash task, which requires precise alignment to a specific area to open the trash bin, DTP demonstrated superior guidance capabilities. While other baseline methods positioned the arm near the target, they often failed to locate the precise opening mechanism, leading to task failure.
In Table III (bottom), we evaluate variations of the training scenario. While training followed the A–B–C sequence, testing extended to longer sequences of up to five subtasks, such as A-B-C-A-C. This design probes long-horizon challenges: for example, repeating Task A after A-B-C occurs in a new context because the bread is already in the drawer, creating situations unseen during training. We created five such long-sequence variants for evaluation. Each value in Table III indicates the number of successful subtasks (out of five) for our method and the baseline, with “Ave. Len.” reporting the average sequence length achieved. In A-B-C-A-C, the baseline fails at the fourth subtask due to missing bread, while our method continues. In A-C-A-B-C, the direct A→C transition introduces both an unseen scenario and initial pose, limiting the baseline to the first task, whereas our method progresses further. Overall, our approach reaches an average length of 4.6, compared to 2.0 for the baseline. This evaluation highlights the impact of accumulated compounding errors in long-horizon tasks: as sequences lengthen and subtask orders vary, accumulated deviations—such as altered object states or unseen robot poses—lead the baseline to fail, while our method remains more robust.
Ave. Suc. | |||||||||||
| ACT [38] | 0.7 | 0.9 | 0.3 | 0.3 | 0.4 | 0.52 | ||||||||||
| BAKU [39] | 0.0 | 0.5 | 0.2 | 0.2 | 0.3 | 0.24 | ||||||||||
| GR1 [31] | 0.7 | 0.7 | 0.2 | 0.4 | 0.4 | 0.48 | ||||||||||
| DTP (ours) | 0.8 | 0.8 | 0.9 | 0.9 | 0.8 | 0.84 | ||||||||||
| ABCAC | ACABC | CABCA | CACAB | BCACA | Ave. Len. | |||||||||||
| GR1 | 3/5 | 1/5 | 4/5 | 2/5 | 0/5 | 2.0 | ||||||||||
| DTP (ours) | 5/5 | 5/5 | 5/5 | 3/5 | 5/5 | 4.6 |
@@
Visualization of Diffusion Trajectory Model. As shown in Fig. 5, we present the overall visualization of the diffusion trajectory generation phase, tested in both the CALVIN environment and real-world scenarios. The visualizations demonstrate that the trajectories generated by our diffusion trajectory prediction closely align with the ground truth. Even when minor deviations occur, the generated trajectories remain consistent with the robotic arm paths dictated by the language instructions.
@@ V CONCLUSION
The limited availability of robot data and compounding errors in IL make it difficult to generalize long-horizon tasks to unseen poses and environments. We introduce a diffusion trajectory-guided framework that leverages RGB-domain diffusion trajectories to improve policy learning in robot manipulation. Our method augments training data through data augmentation or manually crafted labels, producing more accurate trajectories. It has two stages: (i) training a diffusion trajectory model to generate task-relevant trajectories, and (ii) using them to guide the robot’s manipulation policy. On the CALVIN benchmark, our method outperforms state-of-the-art baselines by an average success rate of 25%, and also shows strong improvements using only robot data as well as in real-world experiments.
For future work, we plan to extend our framework to other state-of-the-art policies, as we believe diffusion trajectories can further enhance their effectiveness. Another direction is to obtain trajectory labels using camera intrinsic and extrinsic parameters, which are often missing in open-source datasets [40]. Track-Anything [29] offers strong object-tracking ability for generating labels, while EgoMimic [41] captures detailed 3D hand tracks with Aria glasses that can be projected into 2D particle trajectories. Such hardware also enables large-scale video pretraining for diffusion trajectory tasks.